# VAGUE: Visual Contexts Clarify Ambiguous Expressions

### v2

@@ Abstract
Human communication often relies on visual cues to resolve ambiguity. While humans can intuitively integrate these cues, AI systems often find it challenging to engage in sophisticated multimodal reasoning. We introduce VAGUE, a benchmark evaluating multimodal AI systems’ ability to integrate visual context for intent disambiguation. VAGUE consists of 1.6K ambiguous textual expressions, each paired with an image and multiple-choice interpretations, where the correct answer is only apparent with visual context. The dataset spans both staged, complex (Visual Commonsense Reasoning) and natural, personal (Ego4D) scenes, ensuring diversity. Our experiments reveal that existing multimodal AI models struggle to infer the speaker’s true intent. While performance consistently improves from the introduction of more visual cues, the overall accuracy remains far below human performance, highlighting a critical gap in multimodal reasoning. Analysis of failure cases demonstrates that current models fail to distinguish true intent from superficial correlations in the visual scene, indicating that they perceive images but do not effectively reason with them. We release our code and data at https://github.com/Hazel-Heejeong-Nam/VAGUE.git.
@@ 2.2 Multimodal Implicature Understanding
Implicature and the ambiguity that arises from it naturally emerge in everyday human conversation, requiring pragmatic understanding [36]. Early research on implicature understanding has primarily been conducted in text-only settings [25, 38, 28], with some studies specifically focusing on figurative language and metaphor [20, 4, 37].
However, the ambiguity of standalone text is inherently limited. To address this, recent studies have extended to multiple modalities. One example is multimodal sarcasm understanding (MSU). WITS [18] and MOSES [19] are benchmarks for sarcasm explanation, both providing the speaker’s emotion and voice tone as cues. DocMSU [8] is a document-level benchmark for sarcasm localization and detection. To improve MSU, EDGE [30], a graph-based approach, achieved strong performance on the WITS [18] benchmark. UR-FUNNY [12] is a benchmark for multimodal humor comprehension, incorporating facial expressions and voice tones as in MSU [18, 19] but focusing on humor. Hessel et al. (2023) [13] introduced a benchmark derived from a Cartoon Caption Contest, exploring humor identification and explanation. Baluja et al. (2024) [1] demonstrated that models benefit from multimodal cues in humor understanding. Memes also involve implicature, with multimodal datasets such as MemeCap [14] and MultiBully-Ex [15] proposed for this task.
However, the cues used in multimodal implicature understanding remain simple, primarily appearing in images with a single main object or person, overlooking the importance of interactions between multiple objects and people in real-world scenarios. These limitations underscore the need for more complex cues in implicature understanding, as addressed in VAGUE.
@@ 3.2.1 Directness: Relevance and Solvability
Relevance alone does not guarantee that a prompt is useful. As outlined in Sec. 3.1, the prompt $p_{d}$ must explicitly request an action that the listener can reasonably perform. This introduces solvability as an additional criterion, which requires that the prompt present a clear and actionable problem. A solvable prompt defines a specific issue that can be addressed independently, ensuring that the listener is not left with multiple competing actions to choose from.
@@ 4.2.1 Direct Expressions
After generating $p_{d}$ for all candidate images, human raters evaluate each prompt based on relevance and solvability, assigning scores from 1 to 5. Only those prompts that receive a rating of 4 or 5 are retained for further use. The detailed rating criteria for human verification and an example of the rating process are provided in Fig. H13.
@@ 4.2.2 Indirect Expressions
In the initial step, the model is prompted to generate three distinct candidate options. Each candidate follows the criteria outlined in Section 3.2.2, but with explicit instructions to incorporate different linguistic strategies: sarcasm, humor, and meme/idiomatic expressions, respectively. This approach ensures diversity in the generated responses. In the second step, human annotators evaluate the three candidates and select the one that best aligns with the intended indirectness. The selected prompt is then rated on a scale from 1 to 5 based on more specific criteria. Only those prompts that receive a score of 3 or higher are retained for use in the dataset. The detailed rating criteria for human verification and an example of rating process are provided in Fig. H14
@@ 5 Experiments
We use the following multimodal models in our experiments. The detailed characteristics and descriptions of each model are provided in Appendix C.
@@ 5.1 MLLMs Benefit from Visual Cues
Our first objective is to assess how effectively MLLMs leverage visual cues to resolve ambiguity in utterances. To this end, we systematically control the level of detail in the visual cues provided to the models and measure their accuracy in inferring the speaker’s true intent. Performance is evaluated in both multiple-choice and free-form settings. For clarity, we primarily report multiple-choice accuracy, deferring free-form results to Appendix G.
@@
  • Socratic Models (SMs) [44] use text-only LMs but incorporate short image captions (up to two or three sentences) as additional input. This setup provides minimal visual context, which may be insufficient for accurately inferring intent. Each SM model generated its own image captions and used them in subsequent processing.
@@ Results
The results in Tab. 1 indicate that MLLMs can leverage visual cues, albeit to a limited extent, since SMs and VLMs consistently outperform LMs across all evaluated models. Additionally, more detailed visual input generally improves performance, with VLMs surpassing SMs in most cases, except in the case of proprietary models (GPT-4o and Gemini 1.5 Pro). This exception is further analyzed in Sec. 5.2. Both the VCR and Ego4D subsets exhibit similar performance trends, demonstrating the generalizability of our findings across both staged and real-world scenarios. Finally, the consistently low performance of LMs further reinforces the validity of our dataset as a multimodal benchmark.
@@ 5.2 Analysis on Failure Modes
Here, we examine the ways in which models fail to infer the true intent and how these failure patterns vary with the level of visual cues provided. As shown in Fig. 3, our multiple-choice questions include three distinct types of incorrect answer candidates. We assess the model’s raw visual understanding using the Fake Scene Understanding (FS) and Nonexistent Entity (NE) candidates, which test whether the model can correctly interpret the scene without being misled by fabricated or nonexistent elements. Conversely, the Superficial Understanding (SU) candidate evaluates the model’s reasoning ability, testing whether it can go beyond surface-level perception to infer intent. We provide the full table of failure modes in the MCQ setting in Appendix G
@@ Results
Figure 4 illustrates how frequently each model selects different types of incorrect answers instead of the correct intent. Among the error types, Superficial Understanding (SU) is the most common. This indicates that while models generally succeed in recognizing basic visual details, they often fail to reason deeply about those visual cues to accurately infer the underlying intent of the speaker. However, proprietary models (GPT-4o and Gemini 1.5 Pro) exhibit fewer SU-related errors, indicating stronger reasoning capabilities.
Moreover, stronger visual cues improve accuracy across all models and failure types. This improvement highlights the crucial role of visual conditioning in reducing both vision-based errors (FS and NE) and reasoning-related failures (SU).Notably, proprietary models perform better with captioned inputs (SM) than with raw images (VLM). A closer examination of Fig. 4 reveals that this discrepancy arises from vision-based failures (FS and NE) rather than reasoning-centric failure (SU). This indicates that their captioning ability potentially allows them to obtain more sophisticated information while reducing hallucination in detailed image interpretation.
@@ 5.3 Comparison with human
Table 2: Performance across models and humans on a sampled set of 400 questions. The results are sorted in ascending order of performance. Humans outperform models by a margin of over 30%.
To validate our benchmark and establish an upper bound for performance, we assess human accuracy, highlighting the gap between existing models and human capability. This evaluation follows the multiple-choice setup within the VLM setting, using a subset of 400 samples. As shown in Tab. 2, human performance reaches 94%, demonstrating near-perfect accuracy. Although proprietary models significantly outperform other multimodal systems, a notable performance gap ($\sim$30%) still exists compared to human evaluators. This result underscores the significant gap between AI models’ multimodal reasoning capabilities and human-level understanding when inferring hidden intent, suggesting that visual perception alone, even when accurate, is insufficient without deeper cognitive integration. This performance gap is due to the models’ tendency to rely on surface-level text rather than understanding deeper visual-textual implications. Thus, advancing multimodal reasoning likely requires models to integrate higher-order cognitive processes, like commonsense reasoning and pragmatic understanding, into visual interpretation tasks. Refer to Appendix E for details on the selected subset and human evaluation setup.
@@ 5.4 Chain-of-Thought Experiments
Given the strong reasoning demands of multimodal intent deduction, we further explore the effectiveness of Chain-of-Thought (CoT) prompting [41] in enhancing the reasoning capabilities of MLLMs. The CoT prompt templates, provided in Fig. H21 and  Fig. H22, are designed to explicitly ground the reasoning process, reducing hallucinations. This evaluation is conducted exclusively on proprietary models (GPT-4o and Gemini 1.5 Pro) due to their superior suitability for zero-shot CoT-style prompting.
@@ Results
As shown in Tab. 3, CoT prompting improves performance for raw image inputs (VLM), while showing no clear trend and remaining at a similar level for image caption inputs (SM). One possible explanation for this discrepancy is that CoT primarily enhances reasoning by improving grounding and reducing hallucinations. Since image captions inherently contain fewer hallucinations—albeit at the cost of reduced detail—SMs see little additional benefit from CoT prompting. Additionally, the performance improvements observed with CoT prompting are consistent across different types of false answer candidates, suggesting a generalizable effect in enhancing reasoning quality.
@@ 6 Conclusion
We present VAGUE (Visual Contexts ClArify ambiGUous Expressions), a new benchmark aimed at assessing models’ ability to interpret nuanced communication in complex multimodal scenarios. Our results show that models benefit from visual information when inferring the underlying intention of indirect expressions, as evidenced by their improved performance with increasing levels of visual cues. However, a significant disparity persists between machine capabilities and human-level understanding. To gain deeper insights into model inaccuracies, we design multiple-choice questions that explicitly address failure points, enabling a systematic and quantitative evaluation of the reasons behind performance. The primary challenge identified is the tendency of multimodal models to inadequately integrate visual cues, relying instead on the literal interpretation of textual information. This shortcoming highlights the need for further research, and we anticipate that VAGUE will open up promising avenues for developing systems capable of deeper multimodal reasoning to enhance AI’s ability to engage in human-like interactions.
@@ 7 Limitations
First, there are cultural and linguistic limitations. We incorporate sarcastic, humorous, and idiomatic implicatures in generating indirect expressions. However, since the initial data drafts are created by a model (GPT-4o [29]), they may reflect cultural biases present in its training data. To prevent any potential ethical issues, all human annotators are instructed to remove any content considered problematic or discriminatory during the rating and filtering process. Also, all textual expressions in VAGUE are limited to English. Therefore, we encourage future exploration of indirect expressions across diverse languages and cultures. The second limitation pertains to the dependency of certain meta-information on the quality of the parent dataset [43] and the performance of the models utilized during the dataset generation process. We observe that bounding box annotations from YOLOv11 [17] are occasionally duplicated, leading to a reported number of people higher than actually present. Likewise, the tagging model RAM [46] sometimes misidentified objects. To address this issue, we remove any corrupted instances that could undermine the integrity of the task during the human rating and filtering process.
@@ Appendix A Directness & Indirectness Examples
Figure H3 provides a more concrete explanation with realistic examples of the direct and indirect expressions defined in the main text. Each expression follows two key criteria, and we avoid examples like those in the “Bad” column while prioritizing those in the “Good” cases.
@@ Appendix B VAGUE Benchmark
B.1 Sambles in VAGUE
Figures H5, H6 and H7 illustrate six examples from our benchmark dataset. In the Visual Language Models (VLMs) setting, the model is presented with an image containing person indicator tags, a question, and the speaker’s indirect utterance, as shown in the figures, to answer a multiple-choice question. For reference, we have included the corresponding direct expression below each sample.
@@ B.4 OCR Experiments for Testing Person Tags
[] model response (a) (b) (c) Phi3.5-Vision-Instruct (4B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. LLaVA-Onevision (7B) red red red Qwen2.5-VL-Instruct (7B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. InternVL-2.5-MPO (8B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. Idefics2 (8B) Red. Red. Red. LLaVA-NeXT-vicuna (13B) Person 1 is wearing a red t-shirt. Person 1’s t-shirt is red. Person 5 is wearing a red t-shirt. Ovis2 (16B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. InternVL-2.5-MPO (26B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. GPT-4o Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. Gemini-1.5-Pro Person 1 is wearing a red t-shirt. Person 1 is wearing a red t-shirt. Person 5 is wearing a red t-shirt.

### v3

@@ Abstract
Human communication often relies on visual cues to resolve ambiguity. While humans can intuitively integrate these cues, AI systems often find it challenging to engage in sophisticated multimodal reasoning. We introduce VAGUE, a benchmark evaluating multimodal AI systems’ ability to integrate visual context for intent disambiguation. VAGUE consists of 1.6K ambiguous textual expressions, each paired with an image and multiple-choice interpretations, where the correct answer is only apparent with visual context. The dataset spans both staged, complex (Visual Commonsense Reasoning) and natural, personal (Ego4D) scenes, ensuring diversity. Our experiments reveal that existing multimodal AI models struggle to infer the speaker’s true intent. While performance consistently improves from the introduction of more visual cues, the overall accuracy remains far below human performance, highlighting a critical gap in multimodal reasoning. Analysis of failure cases demonstrates that current models fail to distinguish true intent from superficial correlations in the visual scene, indicating that they perceive images but do not effectively reason with them. We release our code and data at https://hazel-heejeong-nam.github.io/vague/.
footnotetext: Authors contributed equally to this work.
@@ 2.2 Multimodal Implicature Understanding
Implicature and the ambiguity that arises from it naturally emerge in everyday human conversation, requiring pragmatic understanding [38]. Early research on implicature understanding has primarily been conducted in text-only settings [27, 40, 30], with some studies specifically focusing on figurative language and metaphor [21, 4, 39]. However, since the ambiguity of standalone text is inherently limited, recent studies have extended to multiple modalities. One example is multimodal sarcasm understanding (MSU). WITS [19] and MOSES [20] are benchmarks for sarcasm explanation, both providing the speaker’s emotion and voice tone as cues. DocMSU [8] is a document-level benchmark for sarcasm localization and detection. To improve MSU, EDGE [32], a graph-based approach, achieved strong performance. UR-FUNNY [13] is a benchmark for multimodal humor comprehension, incorporating facial expressions and voice tones as in MSU [19, 20]. Hessel et al. (2023) [14] introduced a benchmark derived from a Cartoon Caption Contest, exploring humor identification and explanation. Baluja et al. (2024) [1] demonstrated that models benefit from multimodal cues in humor understanding. Memes also involve implicature, with multimodal datasets such as MemeCap [15] and MultiBully-Ex [16].
However, the cues used in multimodal implicature understanding remain simple, primarily appearing in images with a single main object or person, overlooking the importance of interactions between multiple objects and people in real-world scenarios. These limitations underscore the need for more complex cues, as addressed in VAGUE.
@@ 3.2.1 Directness: Relevance and Solvability
Relevance alone does not guarantee that a prompt is useful. As outlined in Sec. 3.1, the prompt $p_{d}$ must explicitly request an action that the listener can reasonably perform. This introduces solvability, which requires that the prompt present a clear and actionable problem. A solvable prompt defines a specific issue that can be addressed independently, ensuring that the listener is not left with multiple competing actions to choose from.
@@ 4.2.1 Direct Expressions
After generating $p_{d}$ for all candidate images, human raters evaluate each prompt based on relevance and solvability, assigning scores from 1 to 5. Only those prompts that receive a rating of 4 or 5 are retained for further use. The detailed rating criteria for human verification and an example of the rating process are provided in Fig. J13.
@@ 4.2.2 Indirect Expressions
In the initial step, the model is prompted to generate three distinct candidate options. Each candidate follows the criteria outlined in Section 3.2.2, but with explicit instructions to incorporate different linguistic strategies: sarcasm, humor, and meme/idiomatic expressions, respectively. This approach ensures diversity in the generated responses. In the second step, human annotators evaluate the three candidates and select the one that best aligns with the intended indirectness. The selected prompt is then rated on a scale from 1 to 5 based on more specific criteria. Only those prompts that receive a score of 3 or higher are retained for use in the dataset. The detailed rating criteria for human verification and an example of rating process are provided in Fig. J14
@@ 4.2.3 Counterfactual Choices
InternVL-3 (38B) 24.8 47.2 ($\uparrow$ 22.4) 63.6 ($\uparrow$ 38.8) 18.0 47.5 ($\uparrow$ 29.5) 59.8 ($\uparrow$ 41.8)
Qwen2.5-VL-Instruct (72B) 29.6 55.6 ($\uparrow$ 26.0) 74.2 ($\uparrow$ 44.6) 26.8 59.3 ($\uparrow$ 32.5) 69.8 ($\uparrow$ 43.0)
@@ 5 Experiments
We use the following models in our experiments. The detailed descriptions of each model are in Appendix C.
@@ 5.1 MLLMs Benefit from Visual Cues
Our first objective is to assess how effectively MLLMs leverage visual cues to resolve ambiguity in utterances. To this end, we systematically control the level of detail in the visual cues provided to the models and measure their accuracy in inferring the speaker’s true intent. Performance is evaluated in both multiple-choice and free-form settings. For clarity, we primarily report multiple-choice accuracy, deferring free-form results to Appendix I.
@@
  • Socratic Models (SMs) [47] use text-only LMs but incorporate short image captions (up to two or three sentences) as additional input. This short generic caption may lack sufficient detail, which may be insufficient for accurately inferring intent. Each SM model generated its own image captions and used them in subsequent processing.
@@ Results
The results in Tab. 1 indicate that MLLMs can leverage visual cues, albeit to a limited extent, since SMs and VLMs consistently outperform LMs across all evaluated models. Additionally, more detailed visual input generally improves performance, with VLMs surpassing SMs in most cases, except in the case of proprietary models. This exception is further analyzed in Sec. 5.2. Both the VCR and Ego4D subsets exhibit similar performance trends, demonstrating the generalizability of our findings across both staged and real-world scenarios. Finally, the consistently low performance of LMs further reinforces the validity of our dataset as a multimodal benchmark.
@@ 5.2 Analysis on Failure Modes
Here, we examine the ways in which models fail to infer the true intent and how these failure patterns vary with the level of visual cues provided. As shown in Fig. 3, our multiple-choice questions include three distinct types of incorrect answer candidates. We assess the model’s raw visual understanding using the Fake Scene Understanding (FS) and Nonexistent Entity (NE) candidates, which test whether the model can correctly interpret the scene without being misled by fabricated or nonexistent elements. Conversely, the Superficial Understanding (SU) candidate evaluates the model’s reasoning ability, testing whether it can go beyond surface-level perception to infer intent. We provide the full table of failure modes in the MCQ setting in Appendix I
@@ Results
Figure 4 illustrates how frequently each model selects different types of incorrect answers instead of the correct intent. Among the error types, Superficial Understanding (SU) is the most common. This indicates that while models generally succeed in recognizing basic visual details, they often fail to reason deeply about those visual cues to accurately infer the underlying intent of the speaker. However, proprietary models exhibit fewer SU-related errors, indicating stronger reasoning capabilities.
Moreover, stronger visual cues improve accuracy across all models and failure types. This improvement highlights the crucial role of visual conditioning in reducing both vision-based errors (FS and NE) and reasoning-related failures (SU). Notably, proprietary models perform better with captioned inputs (SM) than with raw images (VLM). A closer examination of Fig. 4 reveals that this discrepancy arises from vision-based failures (FS and NE) rather than reasoning-centric failure (SU). This indicates that their captioning ability potentially allows them to obtain more sophisticated information while reducing hallucination in the images. The supporting experiments and explanations are provided in Appendix F.
@@ 5.3 Comparison with human
Table 2: Performance across models and humans on a sampled set of 400 questions. The results show that humans outperform models by a margin of over 20%.
To validate our benchmark and establish an upper bound for performance, we assess human accuracy, highlighting the gap between existing models and human capability. This evaluation follows the multiple-choice setup within the VLM setting, using a subset of 400 samples. As shown in Tab. 2, human performance reaches 94%, demonstrating near-perfect accuracy. Although proprietary models and some large-sized models show a decent performance, a notable performance gap ($\sim$20%) still exists compared to human evaluators. This result underscores the significant gap between AI models’ multimodal reasoning capabilities and human-level understanding when inferring hidden intent, suggesting that visual perception alone, even when accurate, is insufficient without deeper cognitive integration. This performance gap is due to the models’ tendency to rely on surface-level text rather than understanding deeper visual-textual implications. Thus, advancing multimodal reasoning likely requires models to integrate higher-order cognitive processes, like commonsense reasoning and pragmatic understanding, into visual interpretation tasks. Refer to Appendix E for details on the selected subset and human evaluation setup.
@@ 5.4 Chain-of-Thought Experiments
Given the strong reasoning demands of multimodal intent deduction, we further explore the effectiveness of Chain-of-Thought (CoT) prompting [44] in enhancing the reasoning capabilities of MLLMs. The CoT prompt templates, provided in Fig. J21 and  Fig. J22, are designed to explicitly ground the reasoning process, reducing hallucinations. While our main results focus on proprietary models, owing to their superior suitability for zero-shot CoT reasoning, we also present experiments and analyses for open-source models ranging from 4B to 72B in Appendix G.
@@ Results
As shown in Tab. 3, CoT prompting improves performance for raw image inputs (VLM), while showing no clear trend and remaining at a similar level for image caption inputs (SM). One possible explanation for this discrepancy is that CoT primarily enhances reasoning by improving grounding and reducing hallucinations. Since image captions inherently contain fewer hallucinations, SMs see some benefit from CoT prompting albeit at the cost of reduced detail. Additionally, the performance improvements observed with CoT prompting are consistent across different types of false answer candidates, suggesting a generalizable effect in enhancing reasoning quality.
@@ 6 Conclusion
We present VAGUE (Visual Contexts ClArify ambiGUous Expressions), a new benchmark aimed at assessing models’ ability to interpret nuanced communication in complex multimodal scenarios. Our results show that models benefit from visual information when inferring the underlying intention of indirect expressions, as evidenced by their improved performance with increasing levels of visual cues. However, a significant disparity persists between machine capabilities and humans. To gain deeper insights into model inaccuracies, we design multiple-choice questions that explicitly address failure points, enabling a systematic and quantitative evaluation of the reasons behind performance. The primary challenge identified is the tendency of multimodal models to inadequately integrate visual cues, relying instead on the literal interpretation of textual information. This shortcoming highlights the need for further research, and we anticipate that VAGUE will open up promising avenues for developing systems capable of deeper multimodal reasoning to enhance AI’s ability to engage in human-like interactions.
@@ 7 Limitations
First, we acknowledge that there are cultural and linguistic limitations. We incorporate sarcastic, humorous, and idiomatic implicatures in generating indirect expressions. However, since the initial data drafts are created by a model (GPT-4o [31]), they may reflect cultural biases present in its training data. To prevent any potential ethical issues, all human annotators are instructed to remove any content considered problematic or discriminatory during the rating and filtering process. Also, all textual expressions in VAGUE are limited to English. Therefore, we encourage future exploration of indirect expressions across diverse languages and cultures. The second limitation pertains to the dependency of certain meta-information on the quality of the parent dataset and the performance of the models utilized during the dataset generation. We observe that bounding box annotations from YOLOv11 [18] are occasionally duplicated, leading to a reported number of people higher than actually present. Likewise, the tagging model RAM [49] sometimes misidentified objects. Therefore, we remove any corrupted instances that could undermine the integrity of the task during the human rating and filtering process.
@@ Appendix A Directness & Indirectness Examples
Figure J3 provides a more concrete explanation with realistic examples of the direct and indirect expressions defined in the main text. Each expression follows two key criteria, and we avoid examples like those in the “Bad” column while prioritizing those in the “Good” cases.
@@ Appendix B VAGUE Benchmark
B.1 Samples in VAGUE
Figures J5, J6 and J7 illustrate six examples from our benchmark dataset. In the Visual Language Models (VLMs) setting, the model is presented with an image containing person indicator tags, a question, and the speaker’s indirect utterance, as shown in the figures, to answer a multiple-choice question. For reference, we have included the corresponding direct expression below each sample.
@@ B.4 OCR Experiments for Testing Person Tags
[] model response (a) (b) (c) Phi3.5-Vision-Instruct (4B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. LLaVA-Onevision (7B) red red red Qwen2.5-VL-Instruct (7B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. InternVL-2.5-MPO (8B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. Idefics2 (8B) Red. Red. Red. LLaVA-NeXT-vicuna (13B) Person 1 is wearing a red t-shirt. Person 1’s t-shirt is red. Person 5 is wearing a red t-shirt. Ovis2 (16B) Person1 is wearing a red shirt. Person1 is wearing a red shirt. Person5 is wearing a red shirt. InternVL-2.5-MPO (26B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. InternVL-3 (38B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. Qwen2.5-VL-Instruct (72B) Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person5 is wearing a red shirt. GPT-4o Person 1 is wearing a red shirt. Person 1 is wearing a red shirt. Person 5 is wearing a red shirt. Gemini-1.5-Pro Person 1 is wearing a red t-shirt. Person 1 is wearing a red t-shirt. Person 5 is wearing a red t-shirt.